OpenStudy (prakharluv):

Two small conducting spheres of equal radius have charges +10 µC and -20µC respectively and placed at a distance R from reach other. They exprience force F1. If they are brought in contact and separated to the same distance, they exprience force F2. The ratio of F1 to F2 is (a) 1: 8 (b) -8:1 (c) 1:2 (d) -2:1

OpenStudy (anonymous):

Initially, they attract with a force proportional to F1=(+10)(-20) or -200. On contact, they share charge, with 10pC-20pC=-10pC left over, split evenly between the two. then the force will be repulsive F2=(-5)(-5)=+25. Thus, F1/F2=-200/25=-8:1.